# Spectral gap bounds for reversible hybrid Gibbs chains

### v1

@@ 6 Discussion
To obtain a non-trivial lower bound on $1-\|\hat{T}\|_{\Pi}$ using Corollary 10, it is imperative that $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}\leq C$ for all values of $i$ and $y_{i}$, where $C$ is some constant that is strictly less than one. Relaxing this constraint emerges as a clear avenue for prospective research. The comparative outcomes in Corollaries 10, 13, and 16 are useful only if the exact Gibbs sampler in question admits a non-vanishing spectral gap. However, even in the absence of a positive spectral gap, the comparison results based on Dirichlet forms, as expounded in Theorems 9, 12, and 15, may lead to valuable insights into the convergence behavior of the hybrid samplers (Andrieu et al., 2022a). Additionally, exploring the comparison in terms of the conductance (Lawler and Sokal, 1988), the $s$-conductance (Lovász and Simonovits, 1993), or the approximate spectral gap (Atchadé, 2021) could yield further interesting results. Acknowledgment: The first author was partially supported by the National Science Foundation through grant DMS-2112887. The third author was partially supported by the National Science Foundation through grant DMS-2210849 and an Adobe Data Science Research Award. The authors gratefully acknowledge the 2023 IRSA Conference at the University of Minnesota for catalyzing the collaborative efforts among them. Appendix
@@ Appendix B Proof of Theorem 12
Recall the statement of Theorem 12: Theorem 12.Suppose that there exist constants $C_{L},C_{U}\in[-1,1]$ such that
@@ Appendix C Proof of Theorem 15
Before proceeding to the proof, we first give a useful representation of the operator $\hat{S}$. Define a linear operator $Q_{2}:L_{0}^{2}(\Pi)\to L_{0}^{2}(\Pi)$ as follows: for $f\in L_{0}^{2}(\Pi)$,
@@
corresponds to the operator $V_{1}^{*}Q_{2}^{t}V_{1}$ for any positive integer $t$. In particular, $\hat{S}=\hat{S}_{1}=V_{1}^{*}Q_{2}V_{1}$. Another lemma that we need is the following.
@@
Recall the statement of Theorem 15. Theorem 15. Let $\gamma:\mathsf{Y}_{2}\to[0,1]$ be a measurable function such that $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}\leq\gamma(y_{2})$ for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$. Let $t$ be some positive integer, and let $\alpha_{t}\in[0,1]$ be such that
@@
for $M_{1}$-a.e. $y_{1}\in\mathsf{Y}_{1}$. If either (i) $t$ is even, or (ii) $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, then, for $f\in L_{0}^{2}(M_{2})\setminus\{0\}$,
@@
$1-\frac{\langle f,Sf\rangle_{M_{1}}}{\|f\|_{M_{1}}^{2}}-\alpha_{t}\leq 1\frac{\langle f,\hat{S}f\rangle_{M_{1}}^{t}}{\|f\|_{M_{1}}^{2t}}\leq 1$
@@
where we have used (1) and the concentration of $P_{2,y_{2}}(\cdot)$ and $P_{1,y_{1}}(\cdot)$. Assume that $t$ is even, or that $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$. Then, by Lemma 32, $t$ is even, or $Q_{2}$ is positive semi-definite. Note also that $\|V_{1}f\|_{\Pi}^{2}=\|f\|_{M_{1}}^{2}$. By Lemmas 30 and 31,
@@ Appendix D Proof of Proposition 17
Recall the statement of Proposition 17. Proposition 17. In the context of Examples 1 and 3, for $i=1,\dots,n$ and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, there exists an isometric isomorphism $U_{i,y_{i}}:L_{0}^{2}(P_{i,y_{i}})\to L_{0}^{2}(\Pi_{i,y_{i}})$ such that $Q_{i,y_{i}}=U_{i,y_{i}}^{-1}R_{i,y_{i}}U_{i,y_{i}}$. As a consequence, $Q_{i,y_{i}}$ is positive semi-definite if and only if $R_{i,y_{i}}$ is, and $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}=\|R_{i,y_{i}}\|_{\Pi_{i,y_{i}}}$.
@@
Fix $i=1,\dots,n$ and $y_{i}\in\mathsf{Y}_{i}$ such that $\Pi_{i,y_{i}}$ is a probability measure (in case the conditional probability is well-defined only almost surely). For $f\in L_{0}^{2}(P_{i,y_{i}})$, let $U_{i,y_{i}}f$ be the function on $\mathsf{X}_{i}$ such that $U_{i,y_{i}}f(z)=f(x_{1},\dots,x_{n})$ whenever $x_{i}=z\in\mathsf{X}_{i}$ and $x_{-i}=y_{i}$. Then $\Pi_{i,y_{i}}(U_{i,y_{i}}f)=P_{i,y_{i}}f$, and
@@
This implies that $U_{i,y_{i}}$ is a transformation from $L_{0}^{2}(P_{i,y_{i}})$ to $L_{0}^{2}(\Pi_{i,y_{i}})$, and $\|U_{i,y_{i}}f\|_{\Pi_{i,y_{i}}}=\|f\|_{P_{i,y_{i}}}$ for $f\in L_{0}^{2}(P_{i,y_{i}})$. The transformation is clearly linear. To prove that it is an isometric isomorphism, we only need to show that it is bijective. It is injective since if $U_{i,y_{i}}f=0$ for some $f\in L_{0}^{2}(P_{i,y_{i}})$ then $\|f\|_{P_{i,y_{i}}}=\|U_{i,y_{i}}f\|_{\Pi_{i,y_{i}}}=0$. To see that it surjective, note that, for $g\in L_{0}^{2}(\Pi_{i,y_{i}})$, one can identify a function $f\in L_{0}^{2}(P_{i,y_{i}})$ such that $f(x_{1},\dots,x_{n})=g(z)$ whenever $x_{i}=z\in\mathsf{X}_{i}$, which implies that $U_{i,y_{i}}f=g$. It is evident that, for $f\in L_{0}^{2}(P_{i,y_{i}})$ and $z\in\mathsf{X}_{i}$,
@@
Thus, $Q_{i,y_{i}}=U_{i,y_{i}}^{-1}R_{i,y_{i}}U_{i,y_{i}}$. The final claim of the proposition holds since, for $f\in L_{0}^{2}(P_{i,y_{i}})\setminus\{0\}$,
@@ Appendix E Proof of Lemma 25
Recall the statement of the lemma. Lemma 25. Suppose that (A) holds. Let $\Lambda\subset[k]$ and $\Lambda^{\prime}\subset\Lambda$ be such that |Λ|=normal-$|\Lambda|=\ell$ and |Λ|=superscriptnormal-$|\Lambda^{\prime}|=\ell^{\prime}$ (where $\ell>\ell^{\prime}$). Then $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)=P_{\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)$ for $\Pi$-a.e. $x\in\mathsf{X}$.
@@
where $A_{\Lambda}=\{(x,x^{\prime})\in\mathsf{X}^{2}:\,\psi_{\Lambda}(x^{\prime})\neq\psi_{\Lambda}(x)\}$, is measurable by (H) and (A). In conclusion, for $\Pi$-a.e. $x\in\mathsf{X}$, $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}$ is concentrated on $\psi^{-1}_{\Lambda}(\psi_{\Lambda^{\prime}}(x))$. On the other hand, for $A\in\mathcal{B}$,
@@
The third equality follows from (1) with $\Pi$ replaced by $P_{\Lambda,y_{\Lambda}}$ and $P_{i,y_{i}}$ replaced by $P_{\Lambda,y_{\Lambda};\,\Lambda^{\prime},z_{\Lambda^{\prime}}}$. For $z_{\Lambda^{\prime}}\in\mathsf{Y}_{\Lambda^{\prime}}$, let $\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}}$ be the measure $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}$ if $\psi_{\Lambda^{\prime}}(x)=z_{\Lambda^{\prime}}$. Note that $\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}}$ is uniquely defined because $\psi_{\Lambda}(x)=(x_{i})_{i\not\in\Lambda}$ depends on $x$ only through $\psi_{\Lambda^{\prime}}(x)=(x_{i})_{i\not\in\Lambda^{\prime}}$. By previous analyses, $(\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}})_{z_{\Lambda^{\prime}}\in\mathsf{Y}_{\Lambda^{\prime}}}$ is a disintegration of $\Pi$ with respect to $\psi_{\Lambda^{\prime}}$. Because $\mathsf{X}$ and $\mathsf{Y}_{\Lambda}$ are both Polish, the disintegration is unique up to $M_{\Lambda^{\prime}}$ almost sure equivalence (see, e.g., Chang and Pollard, 1997, Theorem 1). In particular, For $M_{\Lambda}$-a.e. $z_{\Lambda^{\prime}}\in\mathsf{Y}_{\Lambda^{\prime}}$, $\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}}=P_{\Lambda^{\prime},z_{\Lambda^{\prime}}}$. This proves the desired result. ∎

### v2

@@ 6 Discussion
To obtain a non-trivial lower bound on $1-\|\hat{T}\|_{\Pi}$ using Corollary 11, it is imperative that $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}\leq C$ for all values of $i$ and $y_{i}$, where $C$ is some constant that is strictly less than one. Relaxing this constraint emerges as a clear avenue for prospective research.
The comparative outcomes in Corollaries 11, 14, and 16 are useful only if the exact Gibbs sampler in question admits a non-vanishing spectral gap. However, even in the absence of a positive spectral gap, the comparison results based on Dirichlet forms, as expounded in Theorems 10, 13, and 15, may lead to valuable insights into the convergence behavior of the hybrid samplers (Andrieu et al., 2022a). Additionally, exploring the comparison in terms of the conductance (Lawler and Sokal, 1988), the $s$-conductance (Lovász and Simonovits, 1993), or the approximate spectral gap (Atchadé, 2021) could yield further interesting results.
Acknowledgment: The first author was partially supported by the National Science Foundation through grant DMS-2112887. The third author was partially supported by the National Science Foundation through grant DMS-2210849 and an Adobe Data Science Research Award. The authors gratefully acknowledge the 2023 IRSA Conference at the University of Minnesota for catalyzing the collaborative efforts among them.
Appendix
@@ Appendix B Proof of Theorem 13
Recall the statement of Theorem 13:
Theorem 13.Suppose that there exist constants $C_{L},C_{U}\in[-1,1]$ such that
@@ Appendix C Proof of Theorem 15
Before proceeding to the proof, we first give a useful representation of the operator $\hat{S}$.
Define a linear operator $Q_{2}:L_{0}^{2}(\Pi)\to L_{0}^{2}(\Pi)$ as follows: for $f\in L_{0}^{2}(\Pi)$,
@@
corresponds to the operator $V_{1}^{*}Q_{2}^{t}V_{1}$ for any positive integer $t$. In particular, $\hat{S}=\hat{S}_{1}=V_{1}^{*}Q_{2}V_{1}$.
Another lemma that we need is the following.
@@
Recall the statement of Theorem 15.
Theorem 15. Let $\gamma:\mathsf{Y}_{2}\to[0,1]$ be a measurable function such that $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}\leq\gamma(y_{2})$ for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$. Let $t$ be some positive integer, and let $\alpha_{t}\in[0,1]$ be such that
@@
for $M_{1}$-a.e. $y_{1}\in\mathsf{Y}_{1}$. If either (i) $t$ is even, or (ii) $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, then, for $f\in L_{0}^{2}(M_{1})\setminus\{0\}$,
@@
where we have used (1) and the concentration of $P_{2,y_{2}}(\cdot)$ and $P_{1,y_{1}}(\cdot)$.
Assume that $t$ is even, or that $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$. Then, by Lemma 32, $t$ is even, or $Q_{2}$ is positive semi-definite. Note also that $\|V_{1}f\|_{\Pi}^{2}=\|f\|_{M_{1}}^{2}$. By Lemmas 30 and 31,
@@ Appendix D Proof of Proposition 17
Recall the statement of Proposition 17.
Proposition 17. In the context of Examples 1 and 3, for $i=1,\dots,n$ and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, there exists an isometric isomorphism $U_{i,y_{i}}:L_{0}^{2}(P_{i,y_{i}})\to L_{0}^{2}(\Pi_{i,y_{i}})$ such that $Q_{i,y_{i}}=U_{i,y_{i}}^{-1}R_{i,y_{i}}U_{i,y_{i}}$. As a consequence, $Q_{i,y_{i}}$ is positive semi-definite if and only if $R_{i,y_{i}}$ is, and $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}=\|R_{i,y_{i}}\|_{\Pi_{i,y_{i}}}$.
@@
Fix $i=1,\dots,n$ and $y_{i}\in\mathsf{Y}_{i}$ such that $\Pi_{i,y_{i}}$ is a probability measure (in case the conditional probability is well-defined only almost surely).
For $f\in L_{0}^{2}(P_{i,y_{i}})$, let $U_{i,y_{i}}f$ be the function on $\mathsf{X}_{i}$ such that $U_{i,y_{i}}f(z)=f(x_{1},\dots,x_{n})$ whenever $x_{i}=z\in\mathsf{X}_{i}$ and $x_{-i}=y_{i}$. Then $\Pi_{i,y_{i}}(U_{i,y_{i}}f)=P_{i,y_{i}}f$, and
@@
This implies that $U_{i,y_{i}}$ is a transformation from $L_{0}^{2}(P_{i,y_{i}})$ to $L_{0}^{2}(\Pi_{i,y_{i}})$, and $\|U_{i,y_{i}}f\|_{\Pi_{i,y_{i}}}=\|f\|_{P_{i,y_{i}}}$ for $f\in L_{0}^{2}(P_{i,y_{i}})$. The transformation is clearly linear. To prove that it is an isometric isomorphism, we only need to show that it is bijective. It is injective since if $U_{i,y_{i}}f=0$ for some $f\in L_{0}^{2}(P_{i,y_{i}})$ then $\|f\|_{P_{i,y_{i}}}=\|U_{i,y_{i}}f\|_{\Pi_{i,y_{i}}}=0$. To see that it surjective, note that, for $g\in L_{0}^{2}(\Pi_{i,y_{i}})$, one can identify a function $f\in L_{0}^{2}(P_{i,y_{i}})$ such that $f(x_{1},\dots,x_{n})=g(z)$ whenever $x_{i}=z\in\mathsf{X}_{i}$, which implies that $U_{i,y_{i}}f=g$.
It is evident that, for $f\in L_{0}^{2}(P_{i,y_{i}})$ and $z\in\mathsf{X}_{i}$,
@@
Thus, $Q_{i,y_{i}}=U_{i,y_{i}}^{-1}R_{i,y_{i}}U_{i,y_{i}}$.
The final claim of the proposition holds since, for $f\in L_{0}^{2}(P_{i,y_{i}})\setminus\{0\}$,
@@ Appendix E Proof of Lemma 25
Recall the statement of the lemma.
Lemma 25. Suppose that (A) holds. Let $\Lambda\subset[k]$ and $\Lambda^{\prime}\subset\Lambda$ be such that $|\Lambda|=\ell$ and $|\Lambda^{\prime}|=\ell^{\prime}$ (where $\ell>\ell^{\prime}$). Then $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)=P_{\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)$ for $\Pi$-a.e. $x\in\mathsf{X}$.
@@
where $A_{\Lambda}=\{(x,x^{\prime})\in\mathsf{X}^{2}:\,\psi_{\Lambda}(x^{\prime})\neq\psi_{\Lambda}(x)\}$, is measurable by (H) and (A). In conclusion, for $\Pi$-a.e. $x\in\mathsf{X}$, $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}$ is concentrated on $\psi^{-1}_{\Lambda}(\psi_{\Lambda^{\prime}}(x))$.
On the other hand, for $A\in\mathcal{B}$,
@@
The third equality follows from (1) with $\Pi$ replaced by $P_{\Lambda,y_{\Lambda}}$ and $P_{i,y_{i}}$ replaced by $P_{\Lambda,y_{\Lambda};\,\Lambda^{\prime},z_{\Lambda^{\prime}}}$.
For $z_{\Lambda^{\prime}}\in\mathsf{Y}_{\Lambda^{\prime}}$, let $\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}}$ be the measure $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}$ if $\psi_{\Lambda^{\prime}}(x)=z_{\Lambda^{\prime}}$. Note that $\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}}$ is uniquely defined because $\psi_{\Lambda}(x)=(x_{i})_{i\not\in\Lambda}$ depends on $x$ only through $\psi_{\Lambda^{\prime}}(x)=(x_{i})_{i\not\in\Lambda^{\prime}}$. By previous analyses, $(\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}})_{z_{\Lambda^{\prime}}\in\mathsf{Y}_{\Lambda^{\prime}}}$ is a disintegration of $\Pi$ with respect to $\psi_{\Lambda^{\prime}}$. Because $\mathsf{X}$ and $\mathsf{Y}_{\Lambda}$ are both Polish, the disintegration is unique up to $M_{\Lambda^{\prime}}$ almost sure equivalence (see, e.g., Chang and Pollard, 1997, Theorem 1). In particular, For $M_{\Lambda}$-a.e. $z_{\Lambda^{\prime}}\in\mathsf{Y}_{\Lambda^{\prime}}$, $\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}}=P_{\Lambda^{\prime},z_{\Lambda^{\prime}}}$. This proves the desired result. ∎